# On the Optimization Dynamics of RLVR: Gradient Gap and Step Size Thresholds

### v2

@@ E.2.3 Proof of Inequality (55)
Supermartingale and Optional Stopping. Using Hoeffding’s lemma [4, Lemma 2.6] on the conditional expectation, we have
Finally, we bound the remaining term using the sub-exponential ($\psi_{1}$-Orlicz) norm of the response length, $T_{\psi_{1}}$ from ˜3. By definition,
$\mathds{E}\big[\exp\{|\vec{\boldsymbol{o}}\!\,|/T_{\psi_{1}}\}\big]\;\leq\;2\,.$
It follows that
Plugging this back into our bound for $\log\mathds{E}[e^{X}]$ gives
$\log\mathds{E}[e^{X}]\;\leq\;G_{\mathrm{p}}^{2}\eta^{2}\cdot T_{\psi_{1}}\,.$
This establishes inequality (57b) and completes the proof.
Appendix F Further Experiments and Details
F.1 REINFORCE on an MAB Problem
@@
Here, we consider a simplified version of the experiment in Section˜5, for a single context/prompt. We run REINFORCE (with an exact value baseline) on a tabular softmax policy class over a bandit problem with $100$ arms, with a randomly chosen best arm, having reward $1$ with all other arms 0. The logits $\theta_{k}$ are initialized as a constant and updated according to the exact gradient update rule:
@@
We construct three plots in Figure˜3, based on the value function $J(k)$, gradient gap $\Delta\mu(k)$, and the cumulative gradient gap $\sum_{k=0}^{K}[\Delta\mu(k)]_{+}\,\cdot\,\eta_{k}$, colored by the step count $k$. Interestingly, the curve $k\mapsto\Delta\mu_{q}(k)$ behaves as a concave quadratic. This is in fact explainable via calculating the alignment gap using our advantage expression (10):
@@
Altogether, our experiment here reinforces the message that increasing cumulative gradient gap corresponds to convergence as in Corollary˜1.
@@
F.2 Further Elaboration on Contextual Bandit Experiments

### v3

@@ H.2.4 Proof Sketch of Lemma 7
By Assumption 2, $|\xi_{t}|\leq G_{\mathrm{p}}\,\eta_{k}$. Using Hoeffding’s lemma (Boucheron et al., 2013, Lemma 2.6) on the conditional MGF and constructing the supermartingale $M_{t}=\exp(2X_{t}-2G_{\mathrm{p}}^{2}\eta_{k}^{2}\cdot t)$, Doob’s optional stopping theorem gives:
$\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot|q)}\big[\exp(2X-2G_{\mathrm{p}}^{2}\eta_{k}^{2}\cdot|\vec{\boldsymbol{o}}\!\,|)\big]\;\leq\;1\,.$
By Cauchy-Schwarz:
$\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot|q)}[e^{X}]\;\leq\;\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot|q)}\big[\exp(2G_{\mathrm{p}}^{2}\eta_{k}^{2}\cdot|\vec{\boldsymbol{o}}\!\,|)\big]^{1/2}\,.$
By Assumption 3, $\||\vec{\boldsymbol{o}}\!\,|\|_{\psi_{1}}\leq T_{\psi_{1}}$ uniformly over all $q$. Under condition (64), standard sub-exponential tail bounds give:
$\log\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot|q)}[e^{X}]\;\leq\;G_{\mathrm{p}}^{2}\,T_{\psi_{1}}\cdot\eta_{k}^{2}\,.$
Crucially, this bound is uniform in $q$ since Assumption 3 assumes $T_{\psi_{1}}$ is a uniform bound. Therefore:
$\log\mathds{E}_{P_{k}}[e^{X}]\;=\;\log\mathds{E}_{q\sim\mathds{P}_{Q}}\big[\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot|q)}[e^{X}]\big]\;\leq\;\log\mathds{E}_{q\sim\mathds{P}_{Q}}\big[\exp(G_{\mathrm{p}}^{2}\,T_{\psi_{1}}\cdot\eta_{k}^{2})\big]\;=\;G_{\mathrm{p}}^{2}\,T_{\psi_{1}}\cdot\eta_{k}^{2}\,.$
Combining steps (i)–(iii) yields (67), and adding the smoothness term completes the proof.
Appendix I Further Experiments and Details
I.1 REINFORCE on an MAB Problem
@@
Here, we consider a simplified version of the experiment in Subsection 6.1, for a single context/prompt. We run REINFORCE (with an exact value baseline) on a tabular softmax policy class over a bandit problem with $100$ arms, with a randomly chosen best arm, having reward $1$ with all other arms 0. The logits $\theta_{k}$ are initialized as a constant and updated according to the exact gradient update rule:
@@
We construct three plots in Figure 4, based on the value function $J(k)$, gradient gap $\Delta\mu(k)$, and the cumulative gradient gap $\sum_{k=0}^{K}[\Delta\mu(k)]_{+}\,\cdot\,\eta_{k}$, colored by the step count $k$. Interestingly, the curve $k\mapsto\Delta\mu_{q}(k)$ behaves as a concave quadratic. This is in fact explainable via calculating the Gap Alignment using our advantage expression (9):
@@
Altogether, our experiment here reinforces the message that increasing cumulative gradient gap corresponds to convergence as in Corollary 3.1.
@@
I.2 Further Elaboration on Contextual Bandit Experiments
@@
I.3 Details on Language Model Experiments
Each of our training routines was performed on a single NVIDIA H200 SXM GPU (with 141GB of VRAM), using the TRL framework (von Werra et al., 2020) for training GRPO without KL regularization.
For training on GSM8K, we performed 3000 training steps with each step involving 4 prompts per batch, 4 responses per prompt in the GRPO estimator, a learning rate of $5\cdot 10^{-6}$, and maximum prompt and completion sequence lengths of $256$.
For training on DAPO-17k, we performed $1051$ training steps with each step involving 4 prompts per batch, 4 responses per prompt, a learning rate of $10^{-6}$, and max sequence lengths of $2048$. For DeepScaleR, we used the same configuration for $733$ training steps.
To compute the validated accuracies, we held out $20\%$ of the questions in each dataset as a holdout test/validation set. Finally, we used a binary reward function on the exact match of the final answer, with validated accuracies in Figure 2 reported as an average over $5$ tries for each step’s model.